Verify that the given functions form the fundamental set of solution of the

differential equation on the given indicated interval.

y"-2y' + 5y = 0

y= c1ex cos 2x + c22ex sin 2x, (-infinity, +infinity)



1
Expert's answer
2021-09-21T03:14:07-0400

The characteristic equation "k^2-2k+5=0" of the differential equation "y''-2y' + 5y = 0" is equivalent to "(k-1)^2=-4," and hence has the roots "k_1=1+2i" and "k_2=1-2i."

Let us show that "e^x\cos 2x" and "e^x\sin 2x" form the fundamental set of solution.

Consider the equality "ae^x\cos 2x+be^x\sin 2x=0." If "x=0," then "ae^0\cos 0+be^0\sin 0=0" implies "a=0." If "x=\frac{\pi}4," then "ae^0\cos \frac{\pi}2+be^0\sin \frac{\pi}2=0" implies "b=0." Therefore, "e^x\cos 2x" and "e^x\sin 2x" are linearly independent, and hence they form the fundamental set of solution.

We conclude that the general solution is "y=C_1e^x\cos 2x+C_2e^x\sin 2x."